Find the midpoint between the points (7, -7) and (-3, -3).

Answer by checkley71(8403)

THE X MID POINT IS (7+3)/2=10/2=5 OR 7-5=2 IS THE X VALUE
THE Y MID POINT IS (-7+3)/2=4/2=2 OR -7+2=-5 IS THE Y VALUE
ANSWER (2,-5) IS THE MID POINT.
THE PROOF IS IN THE PLOTTING OF THESE 2 POINTS, DRAWING A LINE BETWEEN THEM & LOCATING THE MID POINT (2,-5)